Harris Fabrics computes its plantwide predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 39,000 direct labor-hours would be required for the period’s estimated level of production. The company also estimated $583,000 of fixed manufacturing overhead cost for the coming period and variable manufacturing overhead of $3.00 per direct labor-hour. Harris’s actual manufacturing overhead cost for the year was $772,174 and its actual total direct labor was 39,500 hours. Required: Compute the company’s plantwide predetermined overhead rate for the year. (Round your answer to 2 decimal places.)
Answer #1

Total variable overhead estimated=(39000*3)=$117,000

Total overhead estimated=Total variable overhead estimated+Total fixed overhead estimated

=117000+583000=$700,000

Hence predetermined overhead rate=Total overhead estimated/Estimated direct labor hours

=700,000/39000

=$17.95 per labor hour.(Approx).